OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for x in the first equation. 3x + 2y = 7 x - y + 3 = 0 What is the resulting equation? 3x - y - 3 + 2y = 7 3(y - 3) + 2y = 7 3y - 3 + 2y = 7

OpenStudy (anonymous):

from second equation x = y - 3 so plug this into the first

OpenStudy (anonymous):

important - use parentheses around y-3

OpenStudy (anonymous):

x=1/5 y=16/5

OpenStudy (anonymous):

Look at the question properly lilysue, they are asking for the resulting equation. joeywhite's method is correct.

OpenStudy (anonymous):

second on

OpenStudy (anonymous):

x=y-3 3(y-3)+2y=7->y=16/5 x=16/5-3=1/5